# FM-G-CAM: A Holistic Approach for Explainable AI in Computer Vision

### v1

@@ Abstract
Explainability is an aspect of modern AI that is vital for impact and usability in the real world. The main objective of this paper is to emphasise the need to understand the predictions of Computer Vision models, specifically Convolutional Neural Network (CNN) based models. Existing methods of explaining CNN predictions are mostly based on Gradient-weighted Class Activation Maps (Grad-CAM) and solely focus on a single target class. We show that from the point of the target class selection, we make an assumption on the prediction process, hence neglecting a large portion of the predictor CNN model’s thinking process. In this paper, we present an exhaustive methodology called Fused Multi-class Gradient-weighted Class Activation Map (FM-G-CAM) that considers multiple top predicted classes, which provides a holistic explanation of the predictor CNN’s thinking rationale. We also provide a detailed and comprehensive mathematical and algorithmic description of our method. Furthermore, along with a concise comparison of existing methods, we compare FM-G-CAM with Grad-CAM, highlighting its benefits through real-world practical use cases. Finally, we present an open-source Python library with FM-G-CAM implementation to conveniently generate saliency maps for CNN-based model predictions.
@@ 1 Introduction
Despite the strengths of CNNs, they are often considered a black-box approach [9], that is, their predictions are difficult to analyse and interpret. In critical situations where machine learning models are used in the real world, the process that leads to a certain decision may be just as important as the ability of the model itself. The Gradient Class Activation Mapping (Grad-CAM) approach[10] was proposed to visually display the convolutional processes in an easy-to-understand manner. The Grad-CAM algorithm generates a saliency map that explains the areas of a given image that are most important for a classification prediction.
The Grad-CAM saliency map is based on the activations that the CNN model produces when making a prediction and the gradients of them with respect to them. Based on this foundational theory, more advanced studies have been published that introduce improved techniques, such as the use of the second derivative of the resulting activations with respect to the final prediction[11]. However, all of these studies only consider a single class when they produce their saliency maps. This is commonly the top predicted class. In this study, we argue that, except for binary classification, this does not represent the complete rationale of the model for making a given prediction; the class with the highest probability is shown, regardless of the context of other prediction values. This is mainly because it generates the saliency map on a class that may or may not even be the desired one. The saliency map is thus highly dependent on the model output with the highest probability. The difference in the top-1 and top-5 accuracy rates of models [6, 12] trained on ImageNet [13] argues that even the most accurate models are not always able to predict the correct class as the top prediction. In this study, we present FM-G-CAM, an approach to generate saliency maps considering multiple predicted classes, providing a holistic visual explanation on the CNN predictions.
@@
In this paper, we propose a novel approach to produce a saliency map that could provide a unified approach to the visualisation of multiple class predictions; an example of the proposed approach can be observed in Figure 1. Our approach can be configured to consider any number of top classes to consider when generating the saliency map. The paper begins with a comparison of the existing methods to explain CNN predictions. Then, our new approach, FM-G-CAM and the theory behind it are explained mathematically and implemented as a Python package for public download. Following the theory, a list of practical use cases for which our approach could be effectively used is comparatively presented to the existing methods. Finally, the paper concludes with an in-depth discussion of how FM-G-CAM can be widely used to explain CNN predictions, drawbacks of it, and potential avenues to further improve it. The main scientific contributions of this work are as follows:
  • A concise review of the existing saliency map generation techniques. Similarities, differences and drawbacks of them are discussed in the review, highlighting the gaps that exist. We explain the main three techniques in explaining CNN prediction and investigate further into the Activation-based saliency map generation methods.
  • A Holistic approach to explain CNN prediction with example practical use cases. Addressing one of the main drawbacks we highlighted in the review of the existing work, a novel technique, Fused Multi-class Gradient-weighted Activation Map (FM-G-CAM), is introduced. FM-G-CAM is capable of considering multiple predictor classes in generating saliency maps and thereby providing a holistic explanation to CNN-based model predictions. We also provide an exhaustive mathematical and algorithmic explanation for FM-G-CAM, discussing the main concepts behind the novel technique in-depth.
  • A practical usage examples comparative to Grad-CAM, highlighting FM-G-CAM’s benefits. The examples cover general use cases and a potential use case for Medical AI in Chest X-ray diagnosis. We used open-source, state-of-the-art models and datasets for the prediction tasks.
  • A Python package for convenient generation of saliency maps for CNN predictions. We released this Pytorch-oriented library to the Python Package Index (PyPI), allowing the research community to use FM-G-CAM freely and conveniently via a Python library. The library supports different configurations of FM-G-CAM as well as the original Grad-CAM. We plan to maintain and improve this library to serve as a universal inferencing library that is focused on the Explainability of the predictions.
@@ 2.1 Explaining Convolutions
Convolutions Neural Networks were long considered a black box [14] mainly due to the inability to logically explain CNN predictions. Currently, several methods exist to attempt to visually explain the inner workings of a CNN-based model, including CNN-based transformer models [15, 16, 17]. These visualisation methods predominantly use model properties such as, filter weights[18, 19], hidden neurons[20, 21, 22], and inference values which are discussed in the next section.
@@ 2.1.1 Inference values
Perturbation
Blinded grey patches were used by Zeiler et al [23] to identify the areas of the image that have been affecting the prediction results. The variation of the targeted class score is recorded and then used to generate a saliency map visualising the areas that are important for the prediction. Later, this technique was further advanced, replacing the grey square with different pixel value alteration techniques like masking[24].
Although this technique is relatively straightforward, it is very inefficient since each altered patch has to be evaluated by inferencing the patched image through the model. The technique becomes even more inefficient when there is more than one patched/altered area in the image.
Propagation
This technique only requires the image to be propagated once forward and once backwards [25]. The bare gradients calculated while inferencing is used to generate the saliency map. This technique was much faster than pertubation-based techniques. However, the generated saliency maps were occasionally very noisy. Later, more studies were published addressing this noise issue by adding slight modifications to the back-propagation algorithm [26, 27, 28, 29].
The main drawback of this technique is the noisiness of the saliency maps.
Activation
Saliency maps generated using activations of the models while inferencing is commonly called a Class Activation Map (CAM). This was first introduced by Zhou et at. [30] using the activations of the feature maps at the penultimate layer. This was done by identifying the relevance of the activations to the final fully connected output related to the target class. However, this was only possible when the penultimate layer is a Global Average Pooling layer, which was the case back then for most of the typical CNN architectures. Later, Selvaraju et al. [10] presented a way to generate a saliency map with any CNN-based model by getting the CAM of the last convolutional layer and weigting them with the average of their gradients. This is currently the most popular method for generating saliency maps. Grad-CAM was further advanced by using also the second-order derivative in Grad-CAM++ [11]. The idea of Gradient-weighted Activations then created a series of techniques presented in the section 2.2.
Lower resolution of the saliency map is one of the main drawbacks of this technique.
@@ 2.2 CAM-based techniques in summary
Class activations have shown prominent results in explaining CNN predictions mainly due to the information that CNN activations holds regarding the CNN’s decision making process in predictions. Table 2 presents a summary of the major CAM-based saliency map generation methods widely used.
@@
Grad-CAM [10]
Takes the activations of the last convolutional layer and weights them with the average gradients of the same layer with respect to the final fully connected output related to the target class
XGradCAM [31]
Follows a very similar technique to Grad-CAM in weighting activations. However, the calculated gradients were normalised prior to weighting.
HiResCAM [32]
No improvement is the the resolution of the CAM as the name suggests. However, instead of averaging the calculated gradients, in this study the authors multiply the activations element-wise by its corresponding gradient.
Seg-XRes-CAM [33]
A CAM method inspired by HiResCAM for generating saliency maps for segmentation maps, with a primary focus on object detection.
Ablation-CAM [34]
Generates the saliency map in a similar way to performing a ablation test. Sets the prediction activations to zero and measures the deviation in the prediction score. This measured deviation corresponding to the areas of the original image is then used to create the final saliency map.
Score-CAM [35]
Score-CAM removes the dependency on gradients for the generation of the saliency map using a two-stage approach. Instead of gradients, the importance of the activations is determined by a perturbation method similar to that of Ablation-CAM in the second stage.
Eigen-CAM [36]
Differently from most of the other CAM methods, Eigen-CAM uses the first principle component of the feature maps produced following a prediction. This technique can also generate a saliency map without a target class.
Layer-CAM [37]
Very similar to the Grad-CAM’s core techniques but can use an arbitrarily selected convolutional layer of the predictor CNN model.
@@ 2.3 What makes up a good saliency map?
The activation weighting technique affects the accuracy of the saliency map at the very fundamental level. It matters mainly because it is not only the activation that represents the "thinking" process of the CNN, but also the gradients that correspond to them. However, these gradients can be noisy and could cause unrelated activations to be boosted up in the weighting process [9].
Lui et al. [39] present a detailed analysis of different types of quantified colour maps used for data visualisation. Current CAM methods generally use ’viridis’ or ’jet’ colour maps in presenting the saliency map. Our opinion is that the ’viridis’ colourmap is more appropriate than ’jet’ for saliency maps since multiple colours included in the ’jet’ map can make the map more confusing. Colourmap ’viridis’ is also more suitable since we do not need to represent a narrower ranges for intensity in the map.
Generally, the resolution is low in saliency maps and will depend on the properties of the predictor model. If the predictor model has larger kernels the resulting saliency maps will also have larger resolution. However, the kernels tend to become smaller going deeper into the model. Morbidelli et al. [40] present a way to use image augmentation techniques to increase the resolution of the saliency map while having smaller activation maps. This is useful for classification problems with relatively smaller visual targets.
Lastly, we argue that the saliency map must consider multiple classes if and when available. Most of the CAM methods focus only on one targeted class and do not represent the whole context of the CNN model’s thinking process. As an attempt to address this issue, we introduce FM-G-CAM in this work.
@@ 3 Method
The Fused Multi-class Gradient-weighted Class Activation Map (FM-G-CAM) introduced in this study investigates the reasoning behind the model prediction in a broader way compared to the existing methods, thereby aiming to increase explainability. This section begins with a brief discussion of the background that led to the development of FM-G-CAM. Then, the proposed FM-G-CAM algorithm is presented and described.
According to the relevant literature, there is a knowledge gap in XAI for computer vision predictions where more than one class is possible for an input image. Although techniques such as Deep Feature Factorisation [38] work towards this problem, these approaches focus moreso on dissecting an area classified as a single class into multiple classes. The proposed FM-G-CAM approach aims to provide a more holistic explanation of a CNN prediction without relying on one class.
@@ 3.1 Theory
Building on the basic concepts of the Grad-CAM approach[10], FM-G-CAM targets multiple classes instead of a single class which affect the final prediction results. Grad-CAM focuses on a single output at a time, leaving out important information, especially in the case of multi-class classification. FM-G-CAM attempts to overcome this drawback by producing a multicoloured saliency map which highlights the spatial areas of the image corresponding to the top ($N^{*}$) predictions. This can be arbitrarily selected; in this study, the top 3 classes ($K=3$) are selected.
Here ($N=\left\{Prediction\ class\ numbers\right\}$), ($n\in N$) and ($N^{*}$) is defined as below:
@@
In equation 2, ($\alpha_{c}^{n_{k}^{*}}$) represents the importance matrix with respect to the class ($n_{k}^{*}$) and feature channel ($c$) while ($a$) represents the activated outputs of the convolutions. This will then be weighted based on their corresponding activations:
@@
The matrix ($S^{n_{k}^{*}}$) represent filtered saliency map sized ($I\times J$) related to the class ($n_{k}^{*}$). Each element ($S_{i,j}^{n_{k}^{*}}$) is defined as follows:
@@
As mathematically shown in the above equation 4, the filtered saliency map only contains the highest value across the selected ($K$) classes for each element in a 2D matrix ($S^{n_{k}^{*}}$). Other elements are set to zero.
Then, differently from Grad-CAM, L2-normalisation is performed on a concatenated K-dimensional matrix ($S^{{}^{\prime}}$):
@@
As mentioned, ($K$) is set to 3 in this paper for ease of explanation. Deciding a value for ($K$) is discussed in section 3.3.
The L2-normalisation helps to normalise the feature maps across the classes, making the lower activations more pronounced and visible when visually displayed. A comprehensive visual demonstration and comparison between Grad-CAM and FM-G-CAM is done in section 5.
@@ 3.2 Algorithm
In this section, we present the algorithm that realises the mathematical equations explained in Section 3.1. This can be used as a guide when implementing this technique from scratch. Figure 2 visually shows the step-by-step generation of FM-G-CAM from the generation of Gradient-weighted activation maps to the final FM-G-CAM saliency map.
@@
     $W_{C}\leftarrow$ Gradients of the the activations of class $C$ w.r.t to the selected convolutional layer

### v2

@@ Abstract
Explainability is an aspect of modern AI that is vital for impact and usability in the real world. The main objective of this paper is to emphasise the need to understand the predictions of Computer Vision models, specifically Convolutional Neural Network (CNN) based models. Existing methods of explaining CNN predictions are mostly based on Gradient-weighted Class Activation Maps (Grad-CAM) and solely focus on a single target class; from the point of the target class selection, an assumption is made regarding the prediction process, hence neglecting a large portion of the predictor CNN model’s process. In this paper, we present an exhaustive methodology called Fused Multi-class Gradient-weighted Class Activation Map (FM-G-CAM) that considers multiple top predicted classes which provides a holistic explanation of the predictor CNN’s rationale. We also provide a detailed and comprehensive mathematical and algorithmic description of our method. Furthermore, along with a concise comparison of existing methods, we compare FM-G-CAM with Grad-CAM, quantitatively and qualitatively highlighting its benefits through real-world practical use cases. Finally, we present an open-source Python library with FM-G-CAM implementation to conveniently generate saliency maps for CNN-based model predictions.
@@ 1 Introduction
Despite the strengths of CNNs in the field of computer vision, they are often considered a black-box approach [9], that is, their predictions are often difficult to analyse and interpret. In critical situations where machine learning models are used in the real world, the process that leads to a certain decision may be just as important as the ability of the model itself. The Gradient Class Activation Mapping (Grad-CAM) approach[10] was proposed to visually display the convolutional processes in an easy-to-understand manner. The Grad-CAM algorithm generates a saliency map that explains the areas of a given image in a visual manner that are most important for a classification prediction.
The Grad-CAM saliency map is based on the activations that the CNN model produces when making a prediction and their gradients. Based on this foundational theory, more advanced studies have been published that introduce improved techniques, such as the use of the second derivative of the resulting activations with respect to the final prediction[11]. However, all of these studies only consider a single class when producing saliency maps. This is commonly the top predicted class, i.e., that which the model considers the most likely label. We argue that, with the exception of binary classification problems, this does not represent the complete rationale of the model for making a given prediction; the class with the highest probability is shown regardless of the context of other prediction values. This is mainly because it generates the saliency map on a class that may or may not even be the desired one. The saliency map is thus highly dependent on the model output with the highest probability. The difference in the top-1 and top-5 accuracy rates of models [6, 12] trained on ImageNet [13] argues that even the most accurate models are not always able to predict the correct class as the top prediction. In this study, we present FM-G-CAM, an approach to generate saliency maps considering multiple predicted classes, providing a holistic visual explanation of CNN predictions.
@@
In this paper, we propose a novel approach to produce a saliency map that could provide a unified approach to visualising multiple class predictions; an example of the proposed approach can be observed in Figure 1. Our approach can be configured to consider any number of top classes to consider when generating the saliency map. Our paper begins with a comparison of the existing methods to explain CNN predictions. Then, our new approach, FM-G-CAM and the theory behind it are explained mathematically and implemented as a Python package for public download. Following the theory, a list of practical use cases for which our approach could be effectively used is presented in comparison to the existing methods. Finally, the paper concludes with an in-depth discussion of how FM-G-CAM can be widely used to explain CNN predictions, drawbacks of it, and potential avenues to further improve it. The main scientific contributions of our work are as follows:
  • A concise review of the existing saliency map generation techniques. Similarities, differences, and drawbacks of them are discussed in the review, highlighting the gaps that exist. We explain the main three techniques for explaining CNN prediction and further investigate the activation-based saliency map generation methods.
  • A holistic approach to explaining CNN prediction with examples of practical use cases. Addressing one of the main drawbacks we highlighted in the review of the existing work, a novel technique, Fused Multi-class Gradient-weighted Activation Map (FM-G-CAM), is introduced. FM-G-CAM is capable of considering multiple predictor classes in generating saliency maps and thereby providing a holistic explanation to CNN-based model predictions. We also provide an exhaustive mathematical and algorithmic explanation for FM-G-CAM, discussing in-depth the main concepts behind the novel technique.
  • A quantitative evaluation based on a customised test plan repurposing existing saliency map evaluation techniques. The tests uses a pre-trained model and comparatively evaluates FM-G-CAM against Grad-CAM based on more than 80,000 image samples.
  • Practical usage examples similar to Grad-CAM, highlighting FM-G-CAM’s benefits. The examples cover general use cases and a potential use case for Medical AI in Chest X-ray diagnosis. We used open-source, state-of-the-art models and datasets for the prediction tasks.
  • A Python package for convenient generation of saliency maps for CNN predictions. We released this Pytorch-compatible library to the Python Package Index (PyPI), allowing the research community to use FM-G-CAM freely and conveniently via a Python library. The library supports different configurations of FM-G-CAM as well as the original Grad-CAM.
@@ 2.1 Explaining Convolutions
Convolutions Neural Networks were long considered a black box [14] mainly due to the inability to logically explain CNN predictions. Currently, several methods exist to attempt to visually explain the inner workings of a CNN-based model, including CNN-based transformer models [15, 16, 17]. These visualisation methods predominantly use model properties such as, filter weights[18, 19], hidden neurons[20, 21, 22], and inference values, which are discussed in the next section.
@@ 2.1.1 Inference values
Perturbation Blinded grey patches were used by Zeiler et al [23] to identify the areas of the image that have been affecting the prediction results. The variation of the targeted class score is recorded and then used to generate a saliency map visualising the areas that are important for the prediction. Later, this technique was further advanced, replacing the grey square with different pixel value alteration techniques such as masking[24]. Although this technique is relatively straightforward, it is very inefficient since each altered patch has to be evaluated by inferring the patched image through the model. The technique becomes even more inefficient when there is more than one patched/altered area in the image.
Propagation This technique requires only that the image be propagated once forward and once backward [25]. The bare gradients calculated while inferencing is used to generate the saliency map. This technique was much faster than the pertubation-based techniques. However, the generated saliency maps were occasionally very noisy. Later, more studies have been published that address this noise problem by adding slight modifications to the backpropagation algorithm [26, 27, 28, 29]. The main drawback of this technique is the noise within the saliency maps.
Activation Saliency maps generated using activations of the models while inferencing is commonly called a Class Activation Map (CAM). This was first introduced by Zhou et al. [30] using the activations of the feature maps at the penultimate layer. This was done by identifying the relevance of the activations to the final fully connected output related to the target class. However, this was only possible when the penultimate layer is a Global Average Pooling layer, which was the case back then for most of the typical CNN architectures. Later, Selvaraju et al. [10] presented a way to generate a saliency map with any CNN-based model by getting the CAM of the last convolutional layer and weighting them with the average of their gradients. This is currently the most popular method for generating saliency maps. Grad-CAM was further advanced by also using the second-order derivative in Grad-CAM++ [11]. The idea of Gradient-weighted Activations then created a series of techniques presented in the section 2.2. Lower resolution of the saliency map is one of the main drawbacks of this technique.
@@ 2.2 CAM-based techniques in summary
Class activations have shown prominent results in explaining CNN predictions mainly due to the information that CNN activations hold regarding the CNN decision making process in predictions. Table 2 presents a summary of the major CAM-based saliency map generation methods widely used.
@@
Grad-CAM [10] Takes the activations of the last convolutional layer and weights them with the average gradients of the same layer with respect to the final fully connected output related to the target class
XGradCAM [31] Follows a very similar technique to Grad-CAM in weighting activations. However, the calculated gradients were normalised prior to weighting.
HiResCAM [32] No improvement is the the resolution of the CAM as the name suggests. However, instead of averaging the calculated gradients, in this study the authors multiply the activations element-wise by its corresponding gradient.
Seg-XRes-CAM [33] A CAM method inspired by HiResCAM for generating saliency maps for segmentation maps, with a primary focus on object detection.
Ablation-CAM [34] Generates the saliency map in a similar way to performing a ablation test. Sets the prediction activations to zero and measures the deviation in the prediction score. This measured deviation corresponding to the areas of the original image is then used to create the final saliency map.
Score-CAM [35] Score-CAM removes the dependency on gradients for the generation of the saliency map using a two-stage approach. Instead of gradients, the importance of the activations is determined by a perturbation method similar to that of Ablation-CAM in the second stage.
Eigen-CAM [36] Differently from most of the other CAM methods, Eigen-CAM uses the first principle component of the feature maps produced after a prediction. This technique can also generate a saliency map without a target class.
Layer-CAM [37] Very similar to the Grad-CAM’s core techniques but can use an arbitrarily selected convolutional layer of the predictor CNN model.
@@ 2.3 What makes up a good saliency map?
The activation weighting technique affects the accuracy of the saliency map at the very fundamental level. It matters mainly because it is not only the activation that represents the "thinking" process of the CNN, but also the gradients that correspond to them. However, these gradients can be noisy and could cause unrelated activations to be increased in the weighting process [9].
Lui et al. [39] present a detailed analysis of different types of quantified colour maps used for data visualisation. Current CAM methods generally use viridis or jet colour maps in presenting the saliency map. Our opinion is that the ’viridis’ colourmap is more appropriate than ’jet’ for saliency maps, since multiple colours included in the ’jet’ map can make the map more confusing. Colourmap viridis is also more suitable since we do not need to represent narrower ranges for intensity in the map.
Generally, the resolution is low in saliency maps and will depend on the properties of the predictor model. If the predictor model has larger kernels, the resulting saliency maps will also have larger resolution. However, the kernels tend to become smaller going deeper into the model. Morbidelli et al. [40] present a way to use image augmentation techniques to increase the resolution of the saliency map while having smaller activation maps. This is useful for classification problems with relatively smaller visual targets.
Lastly, we argue that the saliency map must consider multiple classes if and when available. Most of the CAM methods focus only on one targeted class and do not represent the whole context of the CNN model’s thinking process. To address this issue, we introduce FM-G-CAM.
@@ 3 Method
We introduce Fused Multi-class Gradient-weighted Class Activation Map (FM-G-CAM) investigating the reasoning behind the model prediction in a broader way compared to the existing methods, thereby aiming to increase explainability. In this section, we describe the background that led to the development of FM-G-CAM followed by a comprehensive explanation of the proposed FM-G-CAM algorithm.
According to the relevant literature, there is a knowledge gap in XAI for computer vision predictions where more than one class is possible for an input image. Although techniques such as Deep Feature Factorisation [38] work towards this problem, these approaches focus more on dissecting an area classified as a single class into multiple classes. The proposed FM-G-CAM approach aims to provide a more holistic explanation of a CNN prediction without relying on one class.
@@ 3.1 Theory
Building on the basic concepts of the Grad-CAM approach[10], FM-G-CAM targets multiple classes instead of a single class, affecting the final prediction results. Grad-CAM focuses on a single output at a time, leaving out important information, especially in the case of multi-class classification. FM-G-CAM attempts to overcome this drawback by producing a multicoloured saliency map that highlights the spatial areas of the image that correspond to the top ($N^{*}$) predictions. This can be arbitrarily selected; we selected the top 5 classes ($K=5$) in this paper. We empirically found that the values ($K=3$) and ($K=4$) work well with the technique.
Here ($N=\left\{Prediction\ class\ numbers\right\}$), ($n\in N$) and ($N^{*}$) is defined as follows:
@@
In equation 2, ($\alpha_{c}^{n_{k}^{*}}$) represents the importance matrix with respect to the class ($n_{k}^{*}$) and the feature channel ($c$) while ($a$) represents the activated outputs of the convolutions. This is then weighted on the basis of their corresponding activations:
@@
The matrix ($S^{n_{k}^{*}}$) represents the filtered saliency map of size ($I\times J$) related to the class ($n_{k}^{*}$). Each element ($S_{i,j}^{n_{k}^{*}}$) is defined as follows:
@@
As mathematically shown in the above equation 4, the filtered saliency map contains only the highest value across the selected ($K$) classes for each element in a 2D matrix ($S^{n_{k}^{*}}$). Other elements are set to zero.
Then, differently from Grad-CAM, the L2-normalisation is performed on a concatenated K-dimensional matrix ($S^{{}^{\prime}}$):
@@
As mentioned above, we set set ($K$) to 5 in this paper for ease of explanation. Deciding a value for ($K$) is discussed in Section 3.3.
L2-normalisation helps normalise feature maps across the classes, making the lower activations more pronounced and visible when visually displayed. We provide a comprehensive visual demonstration and comparison between Grad-CAM and FM-G-CAM in Section 6.
@@ 3.2 Algorithm
In this section, we present the algorithm that realises the mathematical equations explained in Section 3.1. This can be used as a guide to implement this technique from scratch. Figure 2 visually shows the step-by-step generation of FM-G-CAM from the generation of gradient-weighted activation maps to the final FM-G-CAM saliency map.
@@
     $W_{C}\leftarrow$ Gradients of activations of class $C$ w.r.t. the selected convolutional layer